OpenStudy (anonymous):

What is the solution to the system of equations? 4x+y=2 x-y=3

OpenStudy (anonymous):

x = 3 + y substitute into the first one 4(3+y) + y = 2 12 + 4y + y = 2 5y = -10 y = -2

OpenStudy (anonymous):

so 1, -2? @tacosausage